To calculate the probability of a single six-sided die roll coming up with a 1 or 6, the formula would be:

Select one:

a. P(A|B) = P(A) = 1/6

b. P(A) - P(B) = 1/6 - 1/6 = 0

c. P(A) * P(B) = 1/6 * 1/6 = 1/36

d. P(A) + P(B) = 1/6 + 1/6 = 1/3
Answer #1

There are 6 possible outcomes - 1,2,3,4,5,6.

The probability of rolling a 1 P(A) = 1/6.

The probability of rolling a 2 P(B) = 1/6.

=> Probability of rolling a 1 or a 6 = P(A) + P(B) = 1/6 + 1/6 = 1/3.

d. P(A) + P(B) = 1/6 + 1/6 = 1/3.